# A Fisher-Rao gradient flow for entropic mean-field min-max games

### v1

@@ A.2. Existence and uniqueness of the FR flow
Using Assumption 2, (13), (A.7) and (A.8) we obtain
$\log\frac{\nu_{t}(x)}{\pi(x)}\leq\log R_{\nu}+C_{\nu}+\frac{\sigma^{2}}{2}\left(2\log R_{\nu}+\frac{4}{\sigma^{2}}C_{\nu}\right),$
$\log\frac{\mu_{t}(y)}{\rho(y)}\leq\log R_{\mu}+C_{\mu}+\frac{\sigma^{2}}{2}\left(2\log R_{\mu}+\frac{4}{\sigma^{2}}C_{\mu}\right).$
Hence we can choose $R_{1,\nu}\coloneqq 1+\exp\left(\log R_{\nu}+C_{\nu}+\frac{\sigma^{2}}{2}\left(2\log R_{\nu}+\frac{4}{\sigma^{2}}C_{\nu}\right)\right)$ and $R_{1,\mu}\coloneqq 1+\exp\left(\log R_{\mu}+C_{\mu}+\frac{\sigma^{2}}{2}\left(2\log R_{\mu}+\frac{4}{\sigma^{2}}C_{\mu}\right)\right).$ Note $R_{1,\nu},R_{1,\mu}>1$ are conveniently chosen so that $\log R_{1,\nu},\log R_{1,\mu}>0$ in our subsequent calculations. Obtaining a lower bound on $\frac{\nu_{t}(x)}{\pi(x)}$ and $\frac{\mu_{t}(y)}{\rho(y)}$ follows similarly, by using (12) instead of (13).
Step 3: Existence of the gradient flow on $[0,\infty)$. In order to complete our proof, note that the unique solution $(\nu_{t},\mu_{t})_{t\in[0,T]}$ to (A.1) can also be expressed as
$\nu_{t}(x)=\nu_{0}(x)\exp\left(-\int_{0}^{t}\left(\frac{\delta F}{\delta\nu}(\nu_{s},\mu_{s},x)+\frac{\sigma^{2}}{2}\log\left(\frac{\nu_{s}(x)}{\pi(x)}\right)-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\nu_{s}|\pi)\right)\mathrm{d}s\right),$
$\mu_{t}(y)=\mu_{0}(y)\exp\left(\int_{0}^{t}\left(\frac{\delta F}{\delta\mu}(\nu_{s},\mu_{s},y)-\frac{\sigma^{2}}{2}\log\left(\frac{\mu_{s}(y)}{\rho(y)}\right)+\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\mu_{s}|\rho)\right)\mathrm{d}s\right)\,.$
From 2, (A.7), (A.8) and (19), we obtain for any $t\in[0,T]$
(A.9) $\left|\frac{\delta F}{\delta\nu}(\nu_{t},\mu_{t},x)+\frac{\sigma^{2}}{2}\log\left(\frac{\nu_{t}(x)}{\pi(x)}\right)-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\nu_{t}|\pi)\right|\\ \leq 3C_{\nu}+\frac{\sigma^{2}}{2}\left(\max\{|\log r_{1,\nu}|,\log R_{1,\nu}\}+2\log R_{\nu}\right)=:C_{V,\nu},$
@@
This gives $\|\nu_{t}\|_{TV}\leq\|\nu_{0}\|_{TV}e^{C_{V,\nu}t}$ and $\|\mu_{t}\|_{TV}\leq\|\mu_{0}\|_{TV}e^{C_{V,\mu}t}$, and shows that $\nu_{t}$ and $\mu_{t}$ do not explode in any finite time, hence we obtain a global solution $(\nu_{t},\mu_{t})_{t\in[0,\infty)}$. In particular, the bounds in (A.7), (A.8), (19) and (20) hold for all $t\geq 0$. ∎
@@ Appendix B Notation and definitions
In this section we recall some important definitions. Following [3, Definition $5.43$], we start with the notion of differentiability on the space of probability measure that we utilize throughout the paper.
@@
    (B.1) $F(m^{\prime})-F(m)=\int_{0}^{1}\int_{\mathcal{M}}\frac{\delta F}{\delta m}(m+\varepsilon(m^{\prime}-m),x)\left(m^{\prime}-m\right)(\mathrm{d}x)\mathrm{d}\varepsilon.$
@@
If, for fixed $x\in\mathcal{M},$ the map $m\mapsto\frac{\delta F}{\delta m}(m,x)$ satisfies Definition B.1, we say that $F$ admits a second-order flat derivative denoted by $\frac{\delta^{2}F}{\delta m^{2}}.$ Consequently, by Definition B.1, there exists a functional $\frac{\delta^{2}F}{\delta m^{2}}:\mathcal{P}_{p}(\mathcal{M})\times\mathcal{M}\times\mathcal{M}\rightarrow\mathbb{R}$ such that
(B.2) $\frac{\delta F}{\delta m}(m^{\prime},x)-\frac{\delta F}{\delta m}(m,x)=\int_{0}^{1}\int_{\mathcal{M}}\frac{\delta^{2}F}{\delta m^{2}}(\nu+\varepsilon(m^{\prime}-m),x,x^{\prime})\left(m^{\prime}-m\right)(\mathrm{d}x^{\prime})\mathrm{d}\varepsilon.$
Definition B.2 (TV distance between probability measures; [23], Definition $2.4$).

### v2

@@ A.2 Existence and uniqueness of the FR flow
Using Assumption 2, (15), (33) and (34) we obtain
$\log\frac{\nu_{t}(x)}{\pi(x)}\leq 3\log R_{\nu}+\frac{6}{\sigma^{2}}C_{\nu},$
$\log\frac{\mu_{t}(y)}{\rho(y)}\leq 3\log R_{\mu}+\frac{6}{\sigma^{2}}C_{\mu}.$
Hence we can choose $R_{1,\nu}\coloneqq 1+\exp\left(3\log R_{\nu}+\frac{6}{\sigma^{2}}C_{\nu}\right)$ and $R_{1,\mu}\coloneqq 1+\exp\left(3\log R_{\mu}+\frac{6}{\sigma^{2}}C_{\mu}\right).$ Note $R_{1,\nu},R_{1,\mu}>1$ are conveniently chosen so that $\log R_{1,\nu},\log R_{1,\mu}>0$ in our subsequent calculations. Obtaining a lower bound on $\frac{\nu_{t}(x)}{\pi(x)}$ and $\frac{\mu_{t}(y)}{\rho(y)}$ follows similarly, by using (14) instead of (15).
Step 3: Existence of the gradient flow on $[0,\infty)$. In order to complete our proof, note that the unique solution $(\nu_{t},\mu_{t})_{t\in[0,T]}$ to (27) can also be expressed as
$\begin{split}\nu_{t}(x)&=\nu_{0}(x)\exp\left(-\int_{0}^{t}\left(\frac{\delta F}{\delta\nu}(\nu_{s},\mu_{s},x)+\frac{\sigma^{2}}{2}\log\left(\frac{\nu_{s}(x)}{\pi(x)}\right)-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\nu_{s}|\pi)\right)\mathrm{d}s\right),\\ \mu_{t}(y)&=\mu_{0}(y)\exp\left(\int_{0}^{t}\left(\frac{\delta F}{\delta\mu}(\nu_{s},\mu_{s},y)-\frac{\sigma^{2}}{2}\log\left(\frac{\mu_{s}(y)}{\rho(y)}\right)+\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\mu_{s}|\rho)\right)\mathrm{d}s\right)\,.\end{split}$ (35)
Then it follows that the flows $(\nu_{t},\mu_{t})_{t\in[0,T]}$ are continuous and differentiable in time.
From 2, (33), (34) and (21), we obtain for any $t\in[0,T]$
$\left|\frac{\delta F}{\delta\nu}(\nu_{t},\mu_{t},x)+\frac{\sigma^{2}}{2}\log\left(\frac{\nu_{t}(x)}{\pi(x)}\right)-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\nu_{t}|\pi)\right|\\ \leq 3C_{\nu}+\frac{\sigma^{2}}{2}\left(\max\{|\log r_{1,\nu}|,\log R_{1,\nu}\}+2\log R_{\nu}\right)=:C_{V,\nu},$ (36)
@@
This gives $\|\nu_{t}\|_{TV}\leq\|\nu_{0}\|_{TV}e^{C_{V,\nu}t}$ and $\|\mu_{t}\|_{TV}\leq\|\mu_{0}\|_{TV}e^{C_{V,\mu}t}$, and shows that $\nu_{t}$ and $\mu_{t}$ do not explode in any finite time, hence we obtain a global solution $(\nu_{t},\mu_{t})_{t\in[0,\infty)},$ which is continuous and differentiable in time. In particular, the bounds in (33), (34), (21) and (22) hold for all $t>0.$
@@ Appendix B Notation and definitions
In this section we recall some important definitions. Following Carmona & Delarue (2018, Definition $5.43$), we start with the notion of differentiability on the space of probability measure that we utilize throughout the paper.
@@
    $F(m^{\prime})-F(m)=\int_{0}^{1}\int_{\mathcal{M}}\frac{\delta F}{\delta m}(m+\varepsilon(m^{\prime}-m),x)\left(m^{\prime}-m\right)(\mathrm{d}x)\mathrm{d}\varepsilon.$ (37)
@@
If, for any fixed $x\in\mathcal{M},$ the map $m\mapsto\frac{\delta F}{\delta m}(m,x)$ satisfies Definition B.1, we say that $F$ admits a second-order flat derivative denoted by $\frac{\delta^{2}F}{\delta m^{2}}.$ Consequently, by Definition B.1, there exists a functional $\frac{\delta^{2}F}{\delta m^{2}}:\mathcal{P}_{p}(\mathcal{M})\times\mathcal{M}\times\mathcal{M}\rightarrow\mathbb{R}$ such that
$\frac{\delta F}{\delta m}(m^{\prime},x)-\frac{\delta F}{\delta m}(m,x)=\int_{0}^{1}\int_{\mathcal{M}}\frac{\delta^{2}F}{\delta m^{2}}(\nu+\varepsilon(m^{\prime}-m),x,x^{\prime})\left(m^{\prime}-m\right)(\mathrm{d}x^{\prime})\mathrm{d}\varepsilon.$ (38)
Example B.2.
We provide some simple examples of functions $F:\mathcal{P}_{p}(\mathcal{M})\to\mathbb{R}$ and their flat derivatives.
  1. 1.
    Let $F(m)=\int_{\mathcal{M}}f(x)m(\mathrm{d}x)$, where $f:\mathcal{M}\to\mathbb{R}$ is continuous and has $p$-th order growth. Then
    $\frac{\delta F}{\delta m}(m,x)=f(x),\quad\frac{\delta^{2}F}{\delta m^{2}}(m,x)=0.$
  2. 2.
    Let $F(\nu,\mu)=\int_{\mathcal{M}}\int_{\mathcal{M}}f(x,y)\nu(\mathrm{d}x)\mu(\mathrm{d}y).$ Then
    $\frac{\delta F}{\delta\nu}(\nu,\mu,x)=\int_{\mathcal{M}}f(x,y)\mu(\mathrm{d}y),\quad\frac{\delta F}{\delta\mu}(\mu,\nu,y)=\int_{\mathcal{M}}f(x,y)\nu(\mathrm{d}x),\quad\frac{\delta^{2}F}{\delta\nu^{2}}(\nu,\mu,x)=\frac{\delta^{2}F}{\delta\mu^{2}}(\mu,\nu,y)=0.$
  3. 3.
    Let $F(\nu,\mu)=g\left(\int_{\mathcal{M}}\int_{\mathcal{M}}f(x,y)\nu(\mathrm{d}x)\mu(\mathrm{d}y)\right)$ with $f$ continuous and $g$ twice continuously differentiable. Then using the chain rule we obtain
    $\frac{\delta F}{\delta\nu}(\nu,\mu,x)=g^{\prime}\left(\int_{\mathcal{M}}\int_{\mathcal{M}}f(x,y)\nu(\mathrm{d}x)\mu(\mathrm{d}y)\right)\int_{\mathcal{M}}f(x,y)\mu(\mathrm{d}y),$
    $\frac{\delta F}{\delta\mu}(\mu,\nu,y)=g^{\prime}\left(\int_{\mathcal{M}}\int_{\mathcal{M}}f(x,y)\nu(\mathrm{d}x)\mu(\mathrm{d}y)\right)\int_{\mathcal{M}}f(x,y)\nu(\mathrm{d}x),$
    $\frac{\delta^{2}F}{\delta\mu\delta\nu}(\mu,\nu,x,y)=\frac{\delta^{2}F}{\delta\nu\delta\mu}(\nu,\mu,y,x)=g^{\prime}\left(\int_{\mathcal{M}}\int_{\mathcal{M}}f(x,y)\nu(\mathrm{d}x)\mu(\mathrm{d}y)\right)f(x,y)\\ +\int_{\mathcal{M}}f(x,y)\mu(\mathrm{d}y)g^{\prime\prime}\left(\int_{\mathcal{M}}\int_{\mathcal{M}}f(x,y)\nu(\mathrm{d}x)\mu(\mathrm{d}y)\right)\int_{\mathcal{M}}f(x,y)\nu(\mathrm{d}x),$
    $\frac{\delta^{2}F}{\delta\nu^{2}}(\nu,\mu,x,x^{\prime})=g^{\prime\prime}\left(\int_{\mathcal{M}}\int_{\mathcal{M}}f(x,y)\nu(\mathrm{d}x)\mu(\mathrm{d}y)\right)\int_{\mathcal{M}}f(x,y)\mu(\mathrm{d}y)\int_{\mathcal{M}}f(x^{\prime},y)\mu(\mathrm{d}y),$
    $\frac{\delta^{2}F}{\delta\mu^{2}}(\nu,\mu,y,y^{\prime})=g^{\prime\prime}\left(\int_{\mathcal{M}}\int_{\mathcal{M}}f(x,y)\nu(\mathrm{d}x)\mu(\mathrm{d}y)\right)\int_{\mathcal{M}}f(x,y)\nu(\mathrm{d}x)\int_{\mathcal{M}}f(x,y^{\prime})\nu(\mathrm{d}x).$
Definition B.3 (TV distance between probability measures; (Tsybakov, 2008), Definition $2.4$).
@@
Appendix C A formal derivation of the Fisher-Rao gradient flow
In this section, we follow Lu et al. (2019); Yan et al. (2023); Gallouët & Monsaingeon (2017) to formally derive the Fisher-Rao gradient flow of a function $G:\mathcal{P}_{\text{ac}}(\mathcal{M})\to\mathbb{R},$ with $\mathcal{M}\subseteq\mathbb{R}.$ Following Otto (2001), we define the tangent space $T^{\text{FR}}_{\lambda}(\mathcal{M})$ at any $\lambda\in\mathcal{P}_{\text{ac}}(\mathcal{M})$ by
$T^{\text{FR}}_{\lambda}(\mathcal{M})\coloneqq\left\{\text{functions $\xi$ on $\mathcal{M}$ such that $\int_{\mathcal{M}}\xi(x)\mathrm{d}x=0$}\right\}.$
The tangent space $T^{\text{FR}}_{\lambda}(\mathcal{M})$ can also be identified with
$T^{\text{FR}}_{\lambda}(\mathcal{M})\left\{\xi:\xi(x)=\lambda(x)\left(r(x)-\int_{\mathcal{M}}r(x)\lambda(x)\mathrm{d}x\right),\text{ with }r\in L^{2}\left(\mathcal{M};\lambda\right)\right\}.$
On this tangent space, we define the Riemannian metric $g^{\text{FR}}_{\lambda}:T^{\text{FR}}_{\lambda}(\mathcal{M})\times T^{\text{FR}}_{\lambda}(\mathcal{M})\to\mathbb{R}$ at $\lambda$ by
$g^{\text{FR}}_{\lambda}(\xi_{1},\xi_{2})\coloneqq\int_{\mathcal{M}}\frac{\xi_{1}(x)\cdot\xi_{2}(x)}{\lambda(x)}\mathrm{d}x,\quad\text{for }\xi_{1},\xi_{2}\in T^{\text{FR}}_{\lambda}(\mathcal{M}).$ (39)
Then for any $\xi_{i}(x)=\lambda(x)\left(r_{i}(x)-\int_{\mathcal{M}}r_{i}(x)\lambda(x)\mathrm{d}x\right),$ for $i=1,2,$ the metric $g^{\text{FR}}_{\lambda}$ becomes
$g^{\text{FR}}_{\lambda}(\xi_{1},\xi_{2})=\int_{\mathcal{M}}r_{1}(x)r_{2}(x)\lambda(x)\mathrm{d}x-\int_{\mathcal{M}}r_{1}(x)\lambda(x)\mathrm{d}x\int_{\mathcal{M}}r_{2}(x)\lambda(x)\mathrm{d}x.$
With the metric $g^{\text{FR}}_{\lambda}$, the Fisher-Rao distance FR, defined by (3), can be viewed as the geodesic distance on $\left(\mathcal{P}_{\text{ac}}(\mathcal{M}),g^{\text{FR}}_{\lambda}\right).$
Now, we derive the Fisher-Rao gradient flow for a given function $G:\mathcal{P}_{\text{ac}}(\mathcal{M})\to\mathbb{R},$ which we assume is flat differentiable (cf. Definition B.1). Let $(\lambda_{t})_{t\geq 0}$ be a smooth curve starting from $\lambda_{0}=\lambda\in\mathcal{P}_{\text{ac}}(\mathcal{M})$ with arbitrary tangent vector
$\xi(x)=\partial_{t}\lambda_{t}(x)|_{t=0}=\lambda(x)\left(r(x)-\int_{\mathcal{M}}r(x)\lambda(x)\mathrm{d}x\right).$ (40)
The metric gradient $\operatorname{grad}^{\text{FR}}G\in L^{2}(\mathcal{M};\lambda)$ of $G$ with respect to $g^{\text{FR}}_{\lambda}$ is defined by
$\left.\frac{\mathrm{d}}{\mathrm{d}t}G(\lambda_{t})\right|_{t=0}=g^{\text{FR}}_{\lambda}\left(\operatorname{grad}^{\text{FR}}G(\lambda),\xi\right),$ (41)
and the gradient flow of $G$ on $\left(\mathcal{P}_{\text{ac}}(\mathcal{M}),\text{FR}\right)$ is defined by
$\partial_{t}\lambda_{t}(x)=-\operatorname{grad}^{\text{FR}}G(\lambda_{t})(x).$ (42)
Then, by the chain rule, (40) and (39), we obtain
$\left.\frac{\mathrm{d}}{\mathrm{d}t}G(\lambda_{t})\right|_{t=0}=\int_{\mathcal{M}}\frac{\delta G}{\delta\lambda}(\lambda,x)[\partial_{t}\lambda_{t}(x)]|_{t=0}\mathrm{d}x=\int_{\mathcal{M}}\frac{\delta G}{\delta\lambda}(\lambda,x)\xi(x)\mathrm{d}x=g^{\text{FR}}_{\lambda}\left(\frac{\delta G}{\delta\lambda}(\lambda,\cdot)\lambda,\xi\right).$
Since $\xi\in T^{\text{FR}}_{\lambda}(\mathcal{M})$ is arbitrary, it follows from (41) that
$\operatorname{grad}^{\text{FR}}G(\lambda)(x)=\frac{\delta G}{\delta\lambda}(\lambda,x)\lambda(x),$
and consequently by (42),
$\partial_{t}\lambda_{t}(x)=-\frac{\delta G}{\delta\lambda}(\lambda_{t},x)\lambda_{t}(x).$
Appendix D Example: Convergence of discrete-time Fisher-Rao dynamics for a non-interactive min-max game
In this section, we provide an example of a game for which the convergence of the discrete-time Fisher-Rao dynamics to an MNE (measured in NI error) occurs at rate $\mathcal{O}\left(\frac{1}{\eta n}\right)$ in the non-regularized setting, and at rate $\mathcal{O}\left(e^{-\frac{1}{2}\sigma\eta n}\right)$ (to the unique MNE) when regularization is added to the objective function. Here, $\eta>0$ and $\sigma>0$ denote the discretization step-size and regularization parameter, respectively.
To be precise in what follows, we use the terms “discrete-time Fisher-Rao” and “Mirror Descent-Ascent” as synonyms. As it is argued in Wang & Chizat (2023) (at the end of the first bullet point on page 8) a worst-case example which shows convergence of the Fisher-Rao dynamics in discrete time in terms of the NI error at a rate $\mathcal{O}\left(\frac{1}{\eta n}\right)$ can be constructed by following the arguments from Proposition 5.5, Setting II (see the proof of Proposition 5.1 for this specific setting) in Chizat (2022).
Chizat (2022, Proposition 5.5) discusses the case of minimizing a (non-linear) convex function along the iterates of the Mirror Descent algorithm (see Algorithm 1 in (Chizat, 2022)). In order to establish the connection between the setting in Chizat (2022) and the min-max games setting, we note that removing the terms corresponding to the Wasserstein flow in the Conic Particle Mirror-Prox (CP-MP) algorithm in Wang & Chizat (2023) (see the update rule below Lemma $2.1$) and taking $L=1$ retrieves the Mirror Descent-Ascent (MDA) scheme with step-size $\eta>0$ for min-max games in which players update simultaneously. The zero-step-size limit of the MDA scheme is precisely the Fisher-Rao gradient flow.
Following Setting II in the proof of Proposition 5.1 in Chizat (2022), we construct an example for the convergence of the Fisher-Rao dynamics in discrete time in terms of the NI error at a rate $\mathcal{O}\left(\frac{1}{\eta n}\right)$ for the min-max setting in Wang & Chizat (2023) as follows.
First, recall that the min-max game considered in Wang & Chizat (2023) is given by setting $F=\int_{\mathcal{Y}}\int_{\mathcal{X}}f(x,y)\mu(\mathrm{d}x)\nu(\mathrm{d}y)$ and $\sigma=0$ in (1), i.e.,
$\min_{\mu\in\mathcal{P}(\mathcal{X})}\max_{\nu\in\mathcal{P}(\mathcal{Y})}F^{0}(\mu,\nu),\quad\text{ with }F^{0}(\mu,\nu)\coloneqq\int_{\mathcal{Y}}\int_{\mathcal{X}}f(x,y)\mu(\mathrm{d}x)\nu(\mathrm{d}y).$ (43)
Consider $f(x,y)=\Phi(x)-\Phi(y),$ where $\Phi$ is any smooth function such that $\Phi(x)=\frac{1}{2}\|x\|_{2}^{2}$ when $\|x\|_{2}\leq\frac{1}{2}$ and $\Phi(x)\geq\frac{1}{4}$ otherwise. For this choice of $f,$ initial measures $(\mu^{0},\nu^{0})$ supported on the entire space $\mathcal{X}$ and $\mathcal{Y},$ and any step-size $\eta>0,$ the simultaneous MDA update reads
$\begin{cases}\mu^{n+1}=\operatorname*{arg\,min}_{\mu}\{\int\Phi(x)(\mu-\mu^{n})(\mathrm{d}x)+\frac{1}{\eta}\operatorname{D_{KL}}(\mu,\mu^{n})\},\\ \nu^{n+1}=\operatorname*{arg\,max}_{\nu}\{\int(-\Phi(y))(\nu-\nu^{n})(\mathrm{d}y)-\frac{1}{\eta}\operatorname{D_{KL}}(\nu,\nu^{n})\}.\end{cases}$
Hence, we can write $\mu^{n}(x)\propto\mu^{0}(x)\exp(-n\eta\frac{1}{2}\|x\|_{2}^{2})$ and $\nu^{n}(y)\propto\nu^{0}(y)\exp(-n\eta\frac{1}{2}\|y\|_{2}^{2}),$ which are Gaussian distributions of variance $\frac{1}{\eta n}.$ Therefore, for $V^{0}$ given by (43), we can compute
$\operatorname{NI}(\mu^{n},\nu^{n})=\max_{(\mu,\nu)}\left(F^{0}(\nu^{n},\mu)-F^{0}(\nu,\mu^{n})\right)=\int_{\mathcal{Y}}\Phi(y)\nu^{n}(\mathrm{d}y)+\int_{\mathcal{X}}\Phi(x)\mu^{n}(\mathrm{d}x)-\min_{\nu}\int_{\mathcal{Y}}\Phi(y)\nu(\mathrm{d}y)-\min_{\mu}\int_{\mathcal{X}}\Phi(x)\mu(\mathrm{d}x).$
Note that $\nu\mapsto\int\Phi(y)\nu(\mathrm{d}y)$ and $\mu\mapsto\int\Phi(x)\mu(\mathrm{d}x)$ have unique minimizers $\nu^{*}=\delta_{0}$ and $\mu^{*}=\delta_{0},$ respectively, and $\Phi(0)=0.$ Now, assume for example that $(\mu^{0},\nu^{0})$ are uniform distributions on $\mathcal{Y}$ and $\mathcal{X},$ respectively (recall that $\mathcal{X},\mathcal{Y}$ are assumed to be compact in Wang & Chizat (2023)). Then, following the proof of (Chizat, 2022, Proposition 5.5), we obtain
$\int_{\mathcal{Y}}\Phi(y)\nu^{n}(\mathrm{d}y)\approx\frac{1}{2\eta n},$
and analogously
$\int_{\mathcal{X}}\Phi(x)\mu^{n}(\mathrm{d}x)\approx\frac{1}{2\eta n}.$
Hence,
$\operatorname{NI}(\mu^{n},\nu^{n})=\max_{(\mu,\nu)}(F^{0}(\nu^{n},\mu)-F^{0}(\nu,\mu^{n}))=\int\Phi(y)\nu^{n}(\mathrm{d}y)+\int\Phi(x)\mu^{n}(\mathrm{d}x)\approx\frac{1}{\eta n}.$
Now, consider the entropic regularization of (43), i.e.,
$\min_{\mu\in\mathcal{P}(\mathcal{X})}\max_{\nu\in\mathcal{P}(\mathcal{Y})}F^{\sigma}(\mu,\nu),\quad\text{ with }F^{\sigma}(\mu,\nu)\coloneqq\int_{\mathcal{Y}}\int_{\mathcal{X}}f(x,y)\mu(\mathrm{d}x)\nu(\mathrm{d}y)+\sigma\operatorname{D_{KL}}(\mu|\pi)-\sigma\operatorname{D_{KL}}(\nu|\rho),$
where $\sigma>0$ is the regularization parameter and $\pi,\rho$ are defined as in (1). Using the definition of $f,$ a straightforward calculation shows that
$F^{\sigma}(\mu,\nu)=F^{\sigma}_{\pi}(\mu)-F^{\sigma}_{\rho}(\nu),$
where
$F^{\sigma}_{\pi}(\mu)=\int_{\mathcal{X}}\Phi(x)\mu(\mathrm{d}x)+\sigma\operatorname{D_{KL}}(\mu|\pi),$
$F^{\sigma}_{\rho}(\nu)=\int_{\mathcal{Y}}\Phi(y)\nu(\mathrm{d}y)+\sigma\operatorname{D_{KL}}(\nu|\rho).$
In this case, the simultaneous MDA update reads
$\begin{cases}\mu^{n+1}=\operatorname*{arg\,min}_{\mu}\{\int\frac{\delta F^{\sigma}_{\pi}}{\delta\mu}(\mu^{n},x)(\mu-\mu^{n})(\mathrm{d}x)+\frac{1}{\eta}\operatorname{D_{KL}}(\mu|\mu^{n})\},\\ \nu^{n+1}=\operatorname*{arg\,max}_{\nu}\{-\int\frac{\delta F^{\sigma}_{\rho}}{\delta\nu}(\nu^{n},y)(\nu-\nu^{n})(\mathrm{d}y)-\frac{1}{\eta}\operatorname{D_{KL}}(\nu|\nu^{n})\}.\end{cases}$
Note that since the game is non-interactive due to the definition of $f,$ we essentially need to solve two independent optimization problems. Observe that $F^{\sigma}_{\pi}$ and $F^{\sigma}_{\rho}$ are both $\sigma$-strongly convex relative to $\operatorname{D_{KL}}$ in the sense of Lemma 3.1, i.e., for any $\mu^{\prime},\mu\in\mathcal{P}(\mathcal{X}),$ we have
$F^{\sigma}_{\pi}(\mu^{\prime})-F^{\sigma}_{\pi}(\mu)\geq\int\frac{\delta F^{\sigma}_{\pi}}{\delta\mu}(\mu,x)(\mu^{\prime}-\mu)(\mathrm{d}x)+\sigma\operatorname{D_{KL}}(\mu^{\prime}|\mu),$
and analogously for $F^{\sigma}_{\rho}.$ Thus, we can follow the proof of (Aubin-Frankowski et al., 2022, Theorem $4$) with the distinction that the Bregman divergence in Aubin-Frankowski et al. (2022) is actually the KL divergence $\operatorname{D_{KL}}$ in our case. More specifically, in our setup, equation $(39)$ from the proof of (Aubin-Frankowski et al., 2022, Theorem $4$), becomes
$F^{\sigma}_{\pi}(\mu^{n+1})\leq F^{\sigma}_{\pi}(\mu)-\sigma\operatorname{D_{KL}}(\mu|\mu^{n})+\frac{1}{\eta}\operatorname{D_{KL}}(\mu|\mu^{n})-\frac{1}{\eta}\operatorname{D_{KL}}(\mu|\mu^{n+1}),$
for any $\mu\in\mathcal{P}(\mathcal{X}).$ Setting $\mu_{\sigma}^{*}=\operatorname*{arg\,min}_{\mu}F^{\sigma}_{\pi}(\mu)$ in the inequality above, using the fact that $F^{\sigma}_{\pi}(\mu_{\sigma}^{*})\leq F^{\sigma}_{\pi}(\mu^{n+1}),$ and applying the discrete-time Gronwall’s lemma with the assumption $\sigma\eta\leq 1,$ we can show that
$\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu^{n})\leq\exp(-\sigma\eta n)\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu^{0}).$
Analogously, we have
$\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu^{n})\leq\exp(-\sigma\eta n)\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu^{0}).$
Therefore, we obtain
$\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu^{n})+\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu^{n})\leq\exp(-\sigma\eta n)\left(\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu^{0})+\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu^{0})\right).$
Then, as we argue in the proof of Theorem 2.3, we can consequently show that
$\operatorname{NI}(\mu^{n},\nu^{n})\leq 2C_{\sigma}\exp\left(-\frac{1}{2}\sigma\eta n\right)\sqrt{\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu^{0})+\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu^{0})}.$
In conclusion, in the case where there is no interaction between players, adding regularization leads to exponential convergence of the discrete-time Fisher-Rao dynamics to the unique MNE of the game, whereas no regularization guarantees only $\mathcal{O}\left(\frac{1}{\eta n}\right)$ convergence rate to an MNE.